# In the 39th Conference on Neural Information Processing Systems (NeurIPS 2025) Workshop: Embodied World Models for Decision Making

### v2

@@ Abstract
Recent advances in agent development have focused on scaling model size and raw interaction data, mirroring the successes seen in large language models. However, for complex, long-horizon multi-agent tasks such as robotic soccer, this end-to-end approach often fails due to intractable exploration spaces and sparse rewards. This position paper argues that the next frontier in developing embodied world models is not merely increasing the fidelity or size of environments, but scaling their structural complexity through explicit hierarchical scaffolding. We posit that an effective world model for decision-making must model not only the world’s physics but also its task semantics. Drawing from a systematic review of 2024 research in low-resource multi-agent soccer, we identify a clear trend towards integrating symbolic and hierarchical methods, such as Hierarchical Task Networks (HTNs) and Bayesian Strategy Networks (BSNs), with multi-agent reinforcement learning (MARL). These methods decompose complex goals into manageable subgoals, creating an intrinsic curriculum that shapes agent learning. We propose that such structured environments are essential for bridging the gap between simple, reactive behaviors and sophisticated, strategic team play. We further extend this principle, proposing that this scaffolding can be generalized to other complex domains and dynamically generated by Large Language Models (LLMs), which act as generative world models of tasks. By building environments with explicit, composable task layers, we can guide agent exploration more efficiently, generate meaningful learning signals, and ultimately train more capable and general-purpose agents with fewer resources than purely end-to-end approaches.
@@ 1 Introduction: redefining scale for embodied agents
The paradigm of scaling has demonstrably driven profound advancements in artificial intelligence. In the domain of large language models (LLMs), scaling model parameters, dataset size, and computational resources has unlocked emergent capabilities in reasoning, generation, and comprehension (Brown et al., 2020). A parallel trend is now underway in the development of intelligent agents, where the focus is shifting towards scaling interaction. Environments are not passive testbeds, but are themselves a crucial dimension of scale. They provide the rich, dynamic data from which agents learn adaptive behaviors, planning, and long-term decision-making. To advance world models beyond passive prediction and toward active, goal-driven interaction, hinges on the quality and structure of this interaction data.
@@
However, as we push agents towards more complex, autonomous, and general-purpose capabilities, a critical bottleneck emerges, particularly in multi-agent systems, which consistently face challenges of non-stationarity and decentralized decision-making (Ning and Xie, 2024). In domains characterized by long horizons, combinatorial action spaces, and the need for intricate coordination, such as robotic soccer, the brute-force scaling of interaction data via end-to-end reinforcement learning often proves insufficient. Robotic soccer, as a canonical case study for embodied intelligence, encapsulates these challenges, forcing researchers to develop robust solutions for real-time, strategic cooperation under uncertainty (Sammut, 2010). The challenge is not merely one of reward sparsity, but of task sparsity: the sequence of coordinated actions required to achieve a meaningful outcome (e.g., a goal) is so specific and temporally extended that it is almost impossible to discover through random exploration. An agent will not stumble upon a multi-step passing play by chance, regardless of the scale of the simulation. While we ground our analysis in robotic soccer, this challenge of task sparsity is endemic to many complex domains, from collaborative robotic assembly and supply chain logistics to automated scientific discovery and autonomous driving.
@@
This position paper argues for a paradigm shift in how we approach the design of world models and their corresponding environments. We contend that the next crucial frontier is not the scaling of fidelity, size, or agent populations, but the scaling of structural complexity within the environment itself. An embodied world model must capture more than just physical dynamics; it must represent the semantic structure of tasks. We propose that by embedding explicit hierarchical scaffolding into the design of our environments, we can create a structured learning landscape that guides agents toward sophisticated, strategic behavior far more efficiently than flat simulation environments. This scaffolding involves decomposing overarching tasks into a hierarchy of sub-tasks and subgoals, creating an intrinsic curriculum that makes the learning problem tractable.
@@
To substantiate this position, we conduct a focused, systematic review of the 2024 literature on multi-agent robotic soccer, a canonical challenge problem for multi-agent coordination. Our analysis reveals a clear and compelling trend: the most promising recent approaches are moving away from purely end-to-end MARL and towards hybrid models that integrate symbolic planners and hierarchical reasoning. This paper will demonstrate how these methods implicitly or explicitly leverage task decomposition to overcome the limitations of traditional MARL. We will argue that these successes are not isolated algorithmic tricks but rather evidence of a fundamental principle: to learn complex strategies, agents require world models and environments that are structured to teach them.
@@ 2 Related work
Our position builds upon and seeks to synthesize three established pillars of AI research: Hierarchical Reinforcement Learning (HRL), Curriculum Learning, and the emerging use of LLMs as agents and planners. Our contribution is to advocate for externalizing these principles into the environment’s design, creating a more powerful and generalizable learning paradigm.
@@
3 The core challenge: from reward sparsity to task decomposition
The term "sparse rewards" has become a ubiquitous explanation for the difficulties encountered in training agents for long-horizon tasks. In robotic soccer, the ultimate reward, scoring a goal, is an infrequent event, making it a challenging signal for credit assignment across a long sequence of actions. While techniques like reward shaping, which provide denser, more localized feedback for actions like moving towards the ball or possessing it, can be effective (Li et al., 2024b; Baghi, 2024), they often treat the symptoms rather than the underlying disease. The foundational problem is not simply that the goal is a rare event; it is that the path to that goal requires a long, temporally extended, and specific sequence of coordinated sub-tasks. We posit that the core challenge is one of task decomposition.
@@ 3.2 The inefficiency of brute-force scaling
Furthermore, a focus on extreme computational scale can make research inaccessible and sideline promising alternative paradigms. Our methodological review of the 2024 soccer literature was intentionally focused on low-resource approaches that control agents at a high level (e.g., target positions, kick actions) rather than low-level joint angles. This decision was informed by the observation that some of the most complex end-to-end systems, such as those that learn directly from egocentric vision (Tirumala et al., 2024), require a scale of resources that places them beyond the reach of most academic labs. These represent a brute-force approach to a problem that may be better solved with structural priors. The goal should be to find more efficient paths to intelligence, not just the most computationally expensive ones.
@@ 3.3 Methodology of review
To ground our position in empirical evidence, we conducted a systematic survey of academic papers published in 2024 that addressed multi-agent robotic soccer. Our search criteria targeted papers in reputable journals and conference proceedings containing the terms "multi-agent" and "soccer." We applied several filters to align the review with the problem of strategic, low-resource learning:
@@
This curated selection of recent work forms the basis for the analysis presented in the next section, revealing a consistent pattern of leveraging hierarchy and abstraction to tackle the challenge of task decomposition.
4 Evidence from the field: the rise of hierarchical and symbolic methods
@@ 4.1 Explicit task decomposition with symbolic planners
The most compelling evidence for our position comes from the work of Mu et al. (2024), who introduce the HS-MARL framework. This work stands as a direct confirmation that task decomposition is the central challenge. Instead of relying on RL to discover high-level strategy from scratch, HS-MARL integrates a classical AI planner, a Hierarchical Task Network (HTN) planner called pyHIPOP+, into the MARL loop.
@@ 4.3 Abstraction for efficient and interpretable learning
5 A new paradigm for environment design: principles and implications
The evidence points towards a need to fundamentally rethink how we design environments for agent research. The dominant paradigm of creating high-fidelity, flat simulators and challenging agents to learn from scratch is fundamentally inefficient for teaching strategy. We must shift our focus to designing world models and environments that are themselves structured for learning.
5.1 From flat physics simulators to structured task environments
@@
We propose that the next generation of benchmarks should be architected as structured task environments. This means moving beyond physics and providing explicit support for hierarchical task definition and decomposition as a core feature of the environment’s world model itself. The environment should not just be a place where an agent acts, but a place where an agent learns structure.
5.2 Key features of a scaffolded environment
@@
    A Hierarchical Task API: The environment would expose an API for defining tasks and sub-tasks, along with their dependencies and termination conditions. This would be analogous to the HDDL domains used by Mu et al. (2024), but integrated directly into the environment’s state management. For instance, a researcher could define a PassingPlay task composed of the sub-tasks passer_moves_to_ball, receiver_moves_to_open_space, and passer_kicks_to_receiver.
@@
    • Compositional Generalization Score: Measures an agent’s zero-shot or few-shot performance on a novel, complex task composed of sub-tasks it has already mastered individually.
    • Curriculum Efficiency: Quantifies the reduction in sample complexity ($1/N_{samples}$) or wall-clock time required to reach a target performance level in the scaffolded environment compared to a flat baseline.
    • Scaffolding Brittleness Index: Assesses the performance degradation when the provided task hierarchy is intentionally made sub-optimal. This measures an agent’s robustness and ability to learn despite an imperfect scaffold.
5.3 The role of large language models as dynamic planners
While the HTN-based approach in HS-MARL provides a powerful proof-of-concept, its reliance on manually crafted, domain-specific HDDL files represents a significant engineering bottleneck. Recent work has already demonstrated the power of LLMs to decompose high-level instructions into executable plans for embodied agents (Ahn et al., 2022; Wang et al., 2023). We posit that this principle can be adapted to dynamically structure the learning environment itself, transforming static scaffolding into a flexible, language-driven curriculum. The LLM becomes a generative world model of tasks.
In this paradigm, an environment would not have a fixed task hierarchy. Instead, a high-level goal would be specified in natural language (e.g., "Execute a give-and-go play with the forward on the right wing"). An integrated LLM would then perform common-sense reasoning to decompose this instruction into a plausible, ordered sequence of symbolic subgoals: [1. Player A dribbles towards Player B], [2. Player A passes to Player B], [3. Player A runs into open space past defender], [4. Player B returns pass to Player A]. This sequence would be passed to the environment’s task API, dynamically configuring the intrinsic reward function and success conditions for that specific episode. This approach offers several transformative advantages:
@@ 5.4 Implications for agent architectures and Sim2Real transfer
6 Challenges and future directions
@@ 6.1 The scaffolding design problem
A primary challenge is the origin of the hierarchical task structure itself. If the scaffolds must be meticulously hand-crafted by human experts, we have merely shifted the burden from the learning algorithm to the environment designer. This creates a bottleneck and may imbue the agent with the designer’s own biases. The aforementioned use of LLMs as zero-shot planners is a promising mitigation strategy, but it introduces its own challenges, such as ensuring the logical consistency and physical feasibility of LLM-generated plans. Future work could explore methods for agents to learn the task hierarchy itself, perhaps by identifying bottleneck states in their exploration or by abstracting common sub-trajectories into reusable skills.
@@ 6.2 The risk of over-constraining discovery
A critical question is whether providing explicit structure could stifle an agent’s creativity. A scaffold guides an agent along known solution pathways, but it may discourage the discovery of novel, super-human strategies that lie outside the pre-defined hierarchy. This represents a fundamental trade-off between sample efficiency and open-ended exploration. Environments of the future may need to manage this trade-off explicitly, perhaps by employing curriculum strategies that begin with a rigid scaffold and gradually relax constraints as the agent’s competency grows, a process known as fading. Another approach could involve using stochastic planners that introduce structured variation into the task decomposition.

### v3

@@ Abstract
Recent advances in agent development have focused on scaling model size and raw interaction data, mirroring successes in large language models. However, for complex, long-horizon multi-agent tasks such as robotic soccer, this end-to-end approach often fails due to intractable exploration spaces and sparse rewards. We propose that an effective world model for decision-making must model the world’s physics and also its task semantics. A systematic review of 2024 research in low-resource multi-agent soccer reveals a clear trend towards integrating symbolic and hierarchical methods, such as Hierarchical Task Networks (HTNs) and Bayesian Strategy Networks (BSNs), with multi-agent reinforcement learning (MARL). These methods decompose complex goals into manageable subgoals, creating an intrinsic curriculum that shapes agent learning. We formalize this trend into a framework for Hierarchical Task Environments (HTEs), which are essential for bridging the gap between simple, reactive behaviors and sophisticated, strategic team play. Our framework incorporates the use of Large Language Models (LLMs) as generative world models of tasks, capable of dynamically generating this scaffolding. We argue that HTEs provide a mechanism to guide exploration, generate meaningful learning signals, and train agents to internalize hierarchical structure, enabling the development of more capable and general-purpose agents with greater sample efficiency than purely end-to-end approaches.
@@ 1 Introduction: Redefining scale for embodied agents
The paradigm of scaling has demonstrably driven profound advancements in artificial intelligence. In the domain of large language models (LLMs), scaling model parameters, dataset size, and computational resources has unlocked emergent capabilities in reasoning, generation, and comprehension (Brown et al., 2020). A parallel trend is now underway in the development of intelligent agents, where the focus is shifting towards scaling interaction. Environments, rather than passive testbeds, are themselves a crucial dimension of scale. They provide the rich, dynamic data from which agents learn adaptive behaviors, planning, and long-term decision-making. Advancing world models beyond passive prediction and toward active, goal-driven interaction hinges on the quality and structure of this interaction data.
@@
However, as we push agents towards more complex, autonomous, and general-purpose capabilities, a critical bottleneck emerges, particularly in multi-agent systems, which consistently face challenges of non-stationarity and decentralized decision-making (Ning and Xie, 2024). In domains characterized by long horizons, combinatorial action spaces, and the need for intricate coordination, such as robotic soccer, the brute-force scaling of interaction data via end-to-end reinforcement learning often proves insufficient. Robotic soccer, as a canonical case study for embodied intelligence, encapsulates these challenges, forcing researchers to develop robust solutions for real-time, strategic cooperation under uncertainty (Sammut, 2010). The challenge is but of task sparsity: the sequence of coordinated actions required to achieve a meaningful outcome (e.g., a goal) is so specific and temporally extended that it is almost impossible to discover through random exploration. An agent will not stumble upon a multi-step passing play by chance, regardless of the scale of the simulation. While we ground our analysis in robotic soccer, this challenge of task sparsity is endemic to many complex domains, from collaborative robotic assembly and autonomous driving to automated scientific discovery.
@@
This paper argues for a paradigm shift in how we approach the design of world models and their corresponding environments. We argue that the next crucial frontier is not the scaling of fidelity, size, or agent populations, but the scaling of structural complexity within the environment itself. An embodied world model must capture more than just physical dynamics; it must represent the semantic structure of tasks. To address this, we propose a framework for designing Hierarchical Task Environments (HTEs), which embed explicit hierarchical scaffolding into the environment’s world model. This scaffolding involves decomposing overarching tasks into a hierarchy of sub-tasks and subgoals, creating an intrinsic curriculum that makes the learning problem tractable. This scaffolding serves as a powerful curriculum, designed to be gradually faded as agents internalize the task structure, forcing them to learn the very hierarchical reasoning skills that were initially externalized in the environment.
@@
To motivate this framework, we first conduct a systematic review of the 2024 literature on multi-agent robotic soccer. Our analysis reveals a clear and compelling trend: the most promising recent approaches are moving away from purely end-to-end MARL and towards hybrid models that integrate symbolic planners and hierarchical reasoning. This paper will demonstrate how these methods implicitly or explicitly leverage task decomposition to overcome the limitations of traditional MARL. Our contributions are threefold: (1) We identify and analyze the problem of task sparsity in complex MARL domains, arguing it is a more fundamental barrier than reward sparsity. (2) We synthesize a clear trend from the 2024 robotic soccer literature, showing a consistent move towards hybrid, hierarchical methods. (3) We propose a novel framework, Hierarchical Task Environments (HTEs), that externalizes task hierarchy into the environment, and we outline a concrete implementation using LLMs as dynamic task planners. We conclude by discussing new evaluation metrics and the implications of this framework for developing generalizable, embodied agents.
@@ 2 Related work
Our framework builds upon and seeks to synthesize three established pillars of AI research: Hierarchical Reinforcement Learning (HRL), Curriculum Learning, and the emerging use of LLMs as agents and planners. Our contribution is to advocate for externalizing these principles into the environment’s design, creating a more powerful and generalizable learning paradigm.
@@
3 Problem formulation: Task sparsity in multi-agent learning
The term "sparse rewards" has become a ubiquitous explanation for the difficulties encountered in training agents for long-horizon tasks. In robotic soccer, the ultimate reward, scoring a goal, is an infrequent event, making it a challenging signal for credit assignment across a long sequence of actions. While techniques like reward shaping, which provide denser, more localized feedback for actions like moving towards the ball or possessing it, can be effective (Li et al., 2024b; Baghi, 2024), they often treat the symptoms rather than the underlying disease. Beyond simply that the goal is a rare event, the foundational problem is that the path to that goal requires a long, temporally extended, and specific sequence of coordinated sub-tasks. We identify the core challenge as one of task decomposition.
@@ 3.2 The inefficiency of brute-force scaling
Furthermore, a focus on extreme computational scale can make research inaccessible and sideline promising alternative paradigms. Our analysis of the 2024 soccer literature (see Section 4) was intentionally focused on low-resource approaches that control agents at a high level (e.g., target positions, kick actions) rather than low-level joint angles. This decision was informed by the observation that some of the most complex end-to-end systems, such as those that learn directly from egocentric vision (Tirumala et al., 2024), require a scale of resources that places them beyond the reach of most academic labs. These represent a brute-force approach to a problem that may be better solved with structural priors. The goal should be to find more efficient paths to intelligence, instead of the most computationally expensive ones.
4 Analysis of hierarchical methods in multi-agent robotic soccer
@@ 4.1 Methodology of review
To ground our proposal in empirical evidence, we conducted a systematic survey of academic papers published in 2024 that addressed multi-agent robotic soccer. Our search criteria targeted papers in reputable journals and conference proceedings containing the terms "multi-agent" and "soccer." We applied several filters to align the review with the problem of strategic, low-resource learning:
@@
This curated selection of recent work forms the basis for the analysis presented in the next section.
@@ 4.2 Explicit task decomposition with symbolic planners
The most compelling evidence for our proposal comes from the work of Mu et al. (2024), who introduce the HS-MARL framework. This work stands as a direct confirmation that task decomposition is the central challenge. Instead of relying on RL to discover high-level strategy from scratch, HS-MARL integrates a classical AI planner, a Hierarchical Task Network (HTN) planner called pyHIPOP+, into the MARL loop.
@@ 4.4 Abstraction for efficient and interpretable learning
5 HTE: A framework for Hierarchical Task Environments
Based on our analysis, we propose a new framework for environment design to address the task sparsity challenge. The dominant paradigm of creating high-fidelity, flat simulators and challenging agents to learn from scratch is fundamentally inefficient for teaching strategy. We must shift our focus to designing world models and environments that are themselves structured for learning.
5.1 From flat physics simulators to Hierarchical Task Environments
@@
We propose that the next generation of benchmarks should be architected as Hierarchical Task Environments (HTEs). This means moving beyond physics and providing explicit support for hierarchical task definition and decomposition as a core feature of the environment’s world model itself. The environment should be a place where an agent learns structure instead of just be a place where an agent acts.
5.2 Framework components
@@
    A hierarchical task API: The environment would expose an API for defining tasks and sub-tasks, along with their dependencies and termination conditions. This would be analogous to the HDDL domains used by Mu et al. (2024), but integrated directly into the environment’s state management. For instance, a researcher could define a PassingPlay task composed of the sub-tasks passer_moves_to_ball, receiver_moves_to_open_space, and passer_kicks_to_receiver. (See Appendix A for a minimal specification).
@@
    • Compositional generalization score: Measures an agent’s zero-shot or few-shot performance on a novel, complex task composed of sub-tasks it has already mastered individually. The protocol would involve training agents on a set of ’base’ sub-tasks (e.g., dribble, pass) and then testing them on a held-out compositional task (e.g., give-and-go) without further training, comparing their success rate to a flat baseline.
    • Curriculum efficiency: Quantifies the reduction in sample complexity ($1/N_{samples}$) or wall-clock time required to reach a target performance level in the scaffolded environment compared to a flat baseline. This is measured as the ratio of samples-to-convergence ($N_{flat}/N_{scaffolded}$) where the flat baseline is a state-of-the-art MARL agent trained in the same environment with only the terminal goal reward (e.g., scoring).
    • Scaffolding brittleness index: Assesses the performance degradation when the provided task hierarchy is intentionally made sub-optimal. We propose measuring this by comparing the final performance of an agent trained with the full scaffold ($P_{full}$) to one trained with a faded scaffold that has been partially removed ($P_{faded}$), and to an agent trained on a sub-optimal or intentionally wrong hierarchy ($P_{wrong}$). A robust agent would show minimal degradation, indicating it has internalized the strategy rather than simply overfitting to the scaffold.
5.3 Dynamic scaffolding with LLM-based task planners
While the HTN-based approach in HS-MARL provides a powerful proof-of-concept, its reliance on manually crafted, domain-specific HDDL files represents a significant engineering bottleneck. Recent work has already demonstrated the power of LLMs to decompose high-level instructions into executable plans for embodied agents (Ahn et al., 2022; Wang et al., 2023). We propose that this principle can be adapted to dynamically structure the learning environment itself, transforming static scaffolding into a flexible, language-driven curriculum. The LLM becomes a generative world model of tasks.
In this paradigm, an environment would not have a fixed task hierarchy. Instead, a high-level goal would be specified in natural language (e.g., "Execute a give-and-go play with the forward on the right wing"). An integrated LLM would then perform common-sense reasoning to decompose this instruction into a plausible, ordered sequence of symbolic subgoals: [1. Player A dribbles towards Player B], [2. Player A passes to Player B], [3. Player A runs into open space past defender], [4. Player B returns pass to Player A]. This sequence would be passed to the environment’s task API (see Appendix A), dynamically configuring the intrinsic reward function and success conditions for that specific episode. This approach offers several transformative advantages:
@@ 5.3 Dynamic scaffolding with LLM-based task planners
LLM-generated plans introduce challenges of logical consistency and physical feasibility, as noted by Ahn et al. (2022). A robust implementation would require a verification layer to augment the LLM. This blueprint would involve (a) symbolic constraint checking (e.g., verifying that is_possessing(ball) is true before dispatching a pass subgoal) and (b) physical feasibility queries to the world model (e.g., "Is the path for this pass clear?"). If a plan fails verification, the system could re-prompt the LLM with the specific error, or rollback to a previously validated state, creating a robust planning-execution loop.
@@ 5.4 Implications for agent architectures and Sim2Real transfer
6 Discussion and future work
@@ 6.1 The scaffolding design problem
A primary challenge is the origin of the hierarchical task structure itself. If the scaffolds must be meticulously hand-crafted by human experts, we have merely shifted the burden from the learning algorithm to the environment designer. This creates a bottleneck and may imbue the agent with the designer’s own biases. The aforementioned use of LLMs as zero-shot planners is a promising mitigation strategy, but it introduces its own challenges, such as ensuring the logical consistency and physical feasibility of LLM-generated plans. Future work could explore methods for agents to learn the task hierarchy itself, perhaps by identifying bottleneck states in their exploration or by abstracting common sub-trajectories into reusable skills. To mitigate this initial "operator burden," a practical bootstrapping recipe could involve learning from demonstration. An initial library of sub-tasks and options could be learned by observing human expert data, with an LLM then labeling and abstracting these demonstrated trajectories into the symbolic task hierarchy. This human-in-the-loop approach seeds the system with a library of primitives, which can then be composed into new tasks.
@@ 6.2 The risk of over-constraining discovery
A critical concern is whether providing explicit structure could stifle discovery or reduce generality. An agent trained in a scaffolded environment might fail in a flat one, having used the hierarchy as a crutch rather than a learning aid.
We argue that this perspective misinterprets the purpose of the scaffold. The scaffold is a curriculum rather than the final environment. Instead of simply solving the scaffolded task, the goal is to use the scaffold to shape the agent’s internal world model to represent task hierarchies. The training process must therefore include scaffold fading: as the agent masters sub-tasks, the explicit environmental rewards and API-level guidance are gradually removed. This process forces the agent to internalize the hierarchical policy, analogous to how a student, after learning with a tutor, must eventually take the test alone. The fading process itself becomes a key part of the curriculum, training the agent to rely on its own learned, internal task-decomposition model. The proposed Scaffolding Brittleness Index is designed precisely to measure an agent’s success in this internalization and ensure it is not overfitted to the training wheels.
@@ 6.3 The grounding problem
6.4 Generality beyond robotic soccer
While we have grounded our proposal in soccer, this environment-as-scaffold paradigm transfers directly to other complex, long-horizon domains. For instance, in collaborative robotic assembly, the Task API could define a hierarchy of install_component tasks, with subgoals for fetch_part, align_part, and secure_part. In autonomous driving, a high-level instruction like "merge onto the highway" could be dynamically decomposed by the environment into a curriculum of subgoals: match_speed_with_traffic, find_safe_gap, execute_lane_change, and establish_follow_distance. In both cases, the environment provides the structural priors needed to make an intractable, long-horizon problem solvable.
6.5 Interaction with MARL pathologies
A scaffolded environment does not eliminate fundamental MARL challenges such as non-stationarity or partial observability, but it does reframe them. By assigning specific, independent sub-tasks to different agents, the hierarchy can stabilize the learning problem and provide a structured solution to credit assignment. However, it also raises new questions: how should an agent behave if its partner, operating under partial observability, fails at a prerequisite sub-task? Future work could explore how these explicit task dependencies interact with the implicit social dynamics and decentralized nature of MARL.